OpenStudy (anonymous):

I need help in solving this one x2 - 2x - 13 = 0

OpenStudy (anonymous):

\[x = \frac{-b \pm \sqrt{b^2 - 4ac}}{2a}\] for ax^2 +bx +c

OpenStudy (anonymous):

try quadatic formula... i began working it out with the first equation but it went in to decimals and i didn't know how many significant figures you would have wanted

OpenStudy (anonymous):

No decimals, straightforward, just plug the numbers in.

OpenStudy (anonymous):

in quadratic formula?

OpenStudy (anonymous):

U can leave simplified radicals as part of the answer.

OpenStudy (anonymous):

I have no clue on how to do this. It is an assignment for class that I will never use again in my life. Trust me, I am fifty years old and have never had this come up in any work I have done. Therefore I just want to pass this class to move on with the others. Thank you for your help.

OpenStudy (anonymous):

Ok, I believe u:-) The answers (using the above formula) are 1 plus or minus sqrt 14\[1 \pm \sqrt{14}\]

OpenStudy (anonymous):

?